# A Data-Centric Approach To Generate Faithful and High Quality Patient Summaries with Large Language Models

### v1

@@ 5.5 Automatic Hallucination Detection
Using medical entities extracted by MedCat for class-agnostic hallucination detection yielded poor results with an F1-Score of 10.2 for MIMIC and 5.8 for generated summaries (see Table 6). The addition of embeddings (MedCat + Em.) to determine entity equality only resulted in marginal improvements. GPT-4 clearly outperformed this approach on both MIMIC and generated data. The best prompting strategy used a class-aware prompt (detailed in Appendix E). GPT-4 exhibited a low recall in class-aware hallucination detection highlighting the need for improved methods for automatic hallucination detection. Additional results for class-specific recall are presented in Table 14.
@@ 6 Discussion
Hallucinations in patient summaries are diverse and complex. We introduced a rigorous protocol for annotating hallucinations and used a simplified setting with the BHC as limited context. Still, it took medical experts between 12 to 18 minutes to annotate a single summary. For the generated patient summaries, only 59% of all annotations were identified by both annotators, with 56% of them being of the same type. Additionally, 39% of all hallucinations were classified as unsupported word which typically indicates more complex hallucination structures. The annotation of MIMIC summaries yielded similar results, aligning with previous findings on consultation notes that reported medium inter-annotator agreement among clinicians (Moramarco et al., 2022). We attempted to automate hallucination detection based on our annotations. We tested UMLS concepts recognized by MedCat combined with embeddings, which showed very poor results on our datasets. Many annotations had different spans than medical concepts, suggesting that medical entity-based approaches may be insufficient (Cai et al., 2022; Adams et al., 2022). Although GPT-4 demonstrated significantly better performance, we view this only as an initial step. These results emphasize the importance of thorough evaluation with human domain experts and highlight the challenges for reliable hallucination detection in the healthcare domain.
@@
The data used for LLM alignment is crucial for generating faithful and high-quality patient summaries. In a study on medical summarization, Adams et al. (2022) demonstrated that automatic revisions of 45k training examples improved the hallucination rate and quality. In our work, we focused on more sample-efficient LLMs, where 100 human-curated training examples were sufficient. We evaluated Llama 2 and GPT-4 for data-centric hallucination reduction, as well as quantitative and qualitative performance. Consistent with previous studies, we found that quantitative metrics do not correlate well with the hallucination rate (Moramarco et al., 2022; Adams et al., 2023) and quality (Van Veen et al., 2023b). Quantitatively, LED-large outperformed both Llama 2 and GPT-4, but the ranking reversed when assessing quality, including hallucinations (consistency). This behavior can be attributed to the qualitative shortcomings of the MIMIC summaries. LED-large, which was extensively trained, most faithfully reproduced MIMIC, as confirmed by similar qualitative results. Llama 2, fine-tuned on 100 examples, showed moderate alignment to MIMIC, placing it midway in both quantitative and qualitative evaluations. The fine-tuning made Llama 2 more susceptible to data-centric hallucination reduction, which had a strong effect. GPT-4 5-shot, aligned on only five examples, largely retained its pre-trained summarization style and was less susceptible to data-centric hallucination reduction. However, we observed that alignment with only five examples already led to slightly lower relevance, simplification, and length compared to GPT-4 0-shot. Hence, this study emphasizes the importance of high-quality training data for fine-tuning and in-context learning of LLMs.
@@
This work has limitations. Firstly, we only considered a limited context, which led to an overestimation of hallucinations in the MIMIC data. Our hallucination annotation process only took into account spans. Although quality ratings were obtained from experts, it would have been preferable to include the patients’ perspective. Lastly, we only included two LLMs limiting the generalizability of our findings.
@@ 7 Conclusion
LLMs are sensitive to training or prompting examples. In the case of patient summaries, we have demonstrated that data-centric interventions by domain experts can reduce hallucinations and improve the quality. Specifically, GPT-4 showed few hallucinations and omission of key facts while receiving high quality ratings making it a strong model for generating patient summaries. Future work should consider different summary formats and other LLMs. Also, clinical evidence for patient summaries is important.
@@ A.2 Results
We performed a performance analysis on the original MIMIC-IV-Note dataset and MIMIC-IV-Note-DI for prediction of the discharge instructions with LED-large (see Table 8). We can observe that our preprocessing pipeline lead to a significant decrease in performance, suggesting that instances that were easier to predict were removed.
@@
MIMIC-IV-Note-DI ($100,175$ context-summary pairs)
@@
Table 7: Full overview of MIMIC-IV-Note-DI. There are two versions: Full context contains all notes prior to the Discharge Instruction (DI). Short context only has the Brief Hospital Course (BHC). [^10] Sentences and words were determined with nltk. Tokens were determined with the Llama 2 tokenizer.
Figure 5: The preprocessing steps performed on MIMIC-IV-Note to obtain the dataset MIMIC-IV-Note-DI. The goal was obtain diverse and free text discharge instructions (DI) as patient summaries.
@@ A.3 Analyzing 100 Included and Excluded Patient Summaries
Dataset R-1normal-$\uparrow$ R-2normal-$\uparrow$ R-3normal-$\uparrow$ R-4normal-$\uparrow$ R-Lnormal-$\uparrow$ BERTnormal-$\uparrow$ DeBERTnormal-$\uparrow$ SARInormal-$\uparrow$ Words
@@
MIMIC-IV-Note unprocessed $51.17$ $31.28$ $24.06$ $20.08$ $39.85$ $89.10$ $70.48$ $60.62$ $162.49$
MIMIC-IV-Note-DI $44.04$ $17.70$ $9.12$ $5.21$ $29.89$ $88.28$ $64.22$ $46.78$ $82.15$
@@ Appendix B Annotations Statistics
Table 9: Annotation statistics for 100 MIMIC and 100 generated summaries. Annotation were done separately and agreed on together.
@@ Appendix C Parameter Tuning
Model R-1normal-$\uparrow$ R-2normal-$\uparrow$ R-3normal-$\uparrow$ R-4normal-$\uparrow$ R-Lnormal-$\uparrow$ BERTnormal-$\uparrow$ DeBERTnormal-$\uparrow$ SARInormal-$\uparrow$ Words
MIMIC-IV-Note-DI dataset - Short context ($100,175$ ex.)
LED-base (all MIMIC ex.) $43.32$ $17.05$ $8.26$ $4.30$ $29.21$ $87.98$ $63.52$ $46.39$ $74.36$
LED-large (all MIMIC ex.) $43.82$ $17.33$ $8.85$ $4.92$ $29.89$ $88.11$ $64.12$ $46.71$ $76.99$
Llama 2 7B (100 MIMIC ex.) $38.36$ $12.66$ $5.13$ $2.24$ $24.73$ $85.68$ $60.23$ $44.12$ $73.13$
Llama 2 70B (100 MIMIC ex.) $40.58$ $14.31$ $6.09$ $2.74$ $26.19$ $86.30$ $61.89$ $45.16$ $76.90$
GPT-4 5-shot (5 MIMIC ex.) $38.80$ $10.78$ $3.55$ $1.12$ $21.98$ $86.67$ $61.30$ $42.88$ $131.86$
@@
MIMIC-IV-Note-DI dataset - Short context - BHC $\leq$ $4,000$ and DI $\geq$ $600$ characters ($26,178$ ex.)
LED-base (all MIMIC ex.) $42.30$ $14.98$ $7.04$ $3.87$ $26.50$ $86.71$ $60.85$ $44.38$ $117.81$
LED-large (all MIMIC ex.) $46.21$ $17.38$ $8.72$ $5.14$ $28.87$ $87.50$ $63.52$ $45.84$ $117.59$
Llama 2 7B (100 MIMIC ex.) $36.95$ $11.92$ $5.12$ $2.53$ $22.73$ $82.44$ $57.07$ $42.41$ $100.54$
Llama 2 70B (100 MIMIC ex.) $41.82$ $13.63$ $5.77$ $2.66$ $24.83$ $86.43$ $61.34$ $43.86$ $114.08$
GPT-4 5-shot (5 MIMIC ex.) $43.07$ $12.91$ $4.79$ $2.13$ $23.91$ $86.80$ $62.08$ $43.40$ $159.68$
@@
Table 12: All performance results for patient summary generation. We tested all models on the full MIMIC-IV-Note-DI dataset and the subset used for labeling. Parameter or prompt tuning was performed for all models as described in Appendix C.
We performed parameter tuning for the LED (Beltagy et al., 2020) and the Llama 2 (Touvron et al., 2023) models. We used the full MIMIC-IV-Note-DI dataset with short context (BHC) containing 100,175 examples and the subset of 26,178 examples used for labeling. We performed a separate parameter tuning on each dataset. For the LED model we performed full fine-tuning using 80% of the data for training. For Llama 2 we used LoRA for parameter-efficient fine-tuning (Hu et al., 2021) on 100 training examples and loaded the model in 8 bit. For both models, we used 100 examples for validation, and 100 examples for testing. The smaller number of validation and testing examples were chosen to reduce the runtime, however, we still consider them sufficient for representative results. We tracked our experiments with Weights & Biases (Biewald, 2020).
@@ Appendix D Prompt Tuning GPT-4
This finding was supported by the quantitative analysis on ten held-out examples from the MIMIC-IV-Note-DI dataset. The results are shown in table 13. In general, prompt three lead to the best results and additional in-context example did not degrade the performance. Guiding the output with the instructions to start with “You were admitted” also lead to a quantitative improvement. Hence, we chose this prompt for all our GPT-4 summarization experiments.
@@ Appendix E Hallucination Detection with GPT-4
Table 14: Recall for different hallucination labels for hallucination detection on 100 examples from MIMIC-IV-Note-DI and 100 generated summaries using partial matching.
@@
Table 16,17,18 show the prompt used to inform GPT-4 of potential error types and the format for annotating the hallucination spans. We can optionally turn off the hallucination class detection (i.e., generating only <error> rather than <error class="hallucination type">), and we compare the results as class-aware and class-agnostic hallucination detection.
1. Relevance: The rating measures how well the summary captures the key points of the brief hospital course. Consider whether all and only the important aspects are contained in the summary. To this end, please label key facts in the context and the summary. 5 = all key points included 4 = 1 key point missing 3 = 2 key point missing 2 = 3 key point missing 1 = more than 3 key points missing 2. Consistency: The rating measures whether the facts in the summary are consistent with the facts in the original brief hospital course. Consider whether the summary does reproduce all facts accurately and does not make up untrue information. To this end, please label all hallucination in the summary according to the protocol. 5 = no finding 4 = 1-2 minor findings 3 = more than 2 minor or 1 major finding 2 = 2 major findings 1 = more than 2 major findings 3. Simplification: The rating measures whether the summary is written in plain language understandable for a patient. Consider medical terms and abbreviations. To this end, please label medical jargon in the summary. 5 = no medical jargon 4 = 1-2 minor unexplained medical term 3 = more than 2 minor or 1 major medical term 2 = 2 major medical terms 1 = more than 2 major medical terms 4. Fluency: This rating measures the quality of individual sentences, are they well-written and grammatically correct. Consider the quality of individual sentences. 5 = all sentences are well-written and correct 4 = 1-2 sentences have minor errors or poor style 3 = more than 2 sentences with minor errors or poor style or 1 sentence with severe error 2 = 2 sentences with major errors 1 = more than 3 major errors 5. Coherence: The rating measures the quality of all sentences collectively, to the fit together and sound naturally. Consider the quality of the summary as a whole. 5 = the summary as a whole is well-written and clearly structured 4 = the summary is mostly well-written and structured 3 = the summary can be followed and has some structure 2 = the summary is hard to follow and has no clear structure 1 = the summary is very hard to follow and very unstructured
@@ F.1 Labeling Key Facts and Medical Jargon
User
You will be given a doctor’s note and you will need to summarize the patient’s brief hospital course.
Let’s do a practice round.
@@ 3. Incorrect fact
See pages - of instructions.pdf

### v2

@@ 5.5 Automatic Hallucination Detection
Using medical entities extracted by MedCat for class-agnostic hallucination detection performed poorly with an F1-Score of 10.2 for doctor-written (MIMIC) and 5.8 for generated summaries (see Table 5). Adding medical embeddings (MedCat + Em.) only resulted in marginal improvements. GPT-4 clearly outperformed this baseline on both datasets. The best prompting strategy used a class-aware prompt (detailed in Appendix E). GPT-4 exhibited low recall in class-aware hallucination detection highlighting the need for improved methods for automatic hallucination detection. Additional results for class-specific recall are presented in Table 13.
@@ 6 Discussion
Hallucinations in patient summaries are diverse and complex. We introduced a rigorous protocol for annotating hallucinations and considered a simplified experimental setting with the BHC as limited context. Still, it took medical experts between 12 and 18 minutes to annotate a single summary. For the generated patient summaries, only 59% of all annotations were identified by both annotators, with 56% of these being of the same type (see Table 8). Additionally, 39% of all hallucinations were classified as unsupported word which typically indicates more complex hallucination structures. The annotation of doctor-written (MIMIC) summaries yielded similar results, aligning with previous findings on consultation notes that reported medium inter-annotator agreement among clinicians (Moramarco et al., 2022). We attempted to automate hallucination detection based on our annotations. We tested UMLS concepts recognized by MedCat combined with embeddings, which showed very poor results on our datasets. Many annotations spans did not align with medical concepts, suggesting that medical entity-based approaches may be insufficient (Cai et al., 2022; Adams et al., 2022). Although GPT-4 demonstrated significantly better performance, we consider this only as an initial step. These results emphasize the importance of thorough evaluation with human domain experts and highlight the challenges for reliable hallucination detection in the healthcare domain.
@@
The data used for LLM alignment is crucial for generating faithful and high-quality patient summaries. In a study on medical summarization, Adams et al. (2022) demonstrated that automatic revisions of 45,000 training examples improved the hallucination rate and quality. In our work, we focused on more sample-efficient LLMs, where 100 human-curated training examples were sufficient. We evaluated Llama 2 and GPT-4 for data-centric hallucination reduction, as well as for quantitative and qualitative performance. Consistent with previous studies, we found that quantitative metrics do not correlate well with the hallucination rate (Moramarco et al., 2022; Adams et al., 2023) and quality (Van Veen et al., 2024). Quantitatively, LED-large outperformed both Llama 2 and GPT-4, but the ranking reversed when assessing quality, including hallucinations (Consistency). This behavior can be attributed to the qualitative shortcomings of the doctor-written (MIMIC) summaries. LED-large, which was extensively trained, most faithfully reproduced MIMIC, as confirmed by similar qualitative results. Llama 2, fine-tuned on 100 examples, showed moderate alignment to MIMIC, placing it midway in both quantitative and qualitative evaluations. The fine-tuning made Llama 2 more susceptible to data-centric hallucination reduction, which had a strong effect. GPT-4 5-shot, aligned with on only five examples, largely retained its pre-trained summarization style and was less susceptible to data-centric hallucination reduction. However, we observed that alignment with only five examples already led to slightly lower Relevance, Simplification, and length compared to GPT-4 without any examples (0-shot). Hence, this study emphasizes the importance of high-quality training data for fine-tuning and in-context learning of LLMs.
@@
This work has limitations. We relied on a single dataset for medical notes and focused on patient summaries, limiting the generalizability of our findings. We only considered the reduced context of the BHC, which likely led to an overestimation of hallucinations in doctor-written summaries. The labeling of hallucinations was conducted by only two annotators, leading to variability in the labels. The qualitative ratings were obtained from medical experts. Instead, it would have been preferable to include the patients’ perspectives. Lastly, we included only two LLMs in our data-centric hallucination-reduction experiments and qualitative evaluation due to the expensive annotation required by medical experts.
@@ 7 Conclusion
LLMs are sensitive to fine-tuning or prompting examples. We have demonstrated that data-centric interventions by domain experts can reduce hallucinations and improve the quality of generation. Specifically, GPT-4 showed few hallucinations and omitted key facts while receiving high-quality ratings making it a strong model for generating patient summaries. We have developed a protocol for labeling unsupported evidence in medical texts and have released two annotated datasets consisting of 100 doctor-written and 100 LLM-generated patient summaries. Future work should explore different summary formats and the application of other LLMs. The provided datasets could serve as a valuable resource for the advancement and assessment of automatic hallucination metrics. A crucial direction for future research is the collection of more clinical evidence for useful interventions to enhance patient health literacy.
@@ A.2 Results
We performed a performance analysis on the original MIMIC-IV-Note dataset and MIMIC-IV-Note-Ext-DI for prediction of the discharge instructions with LED-large (see Table 7). We can observe that our preprocessing pipeline lead to a significant decrease in performance, suggesting that instances that were easier to predict were removed.
@@
MIMIC-IV-Note-Ext-DI ($100,175$ context-summary ex.)
@@
Table 6: Overview of MIMIC-IV-Note-Ext-DI with the full context and MIMIC-IV-Note-Ext-DI-BHC with the Brief Hospital Course (BHC) as context. The discharge summaries (DI serve as patient summaries for both datasets. [^6] Sentences and words were determined with nltk. Tokens were determined with the Llama 2 tokenizer.
Figure 5: The preprocessing steps performed on MIMIC-IV-Note to obtain the datasets MIMIC-IV-Note-Ext-DI(-BHC). The goal was obtain diverse and free text discharge instructions (DI) as patient summaries.
@@ A.3 Analyzing 100 Included and Excluded Patient Summaries
[…]
@@
MIMIC-IV-Note (unprocessed) $51.17$ $31.28$ $24.06$ $20.08$ $39.85$ $89.10$ $70.48$ $60.62$ $162.49$
MIMIC-IV-Note-Ext-DI $44.04$ $17.70$ $9.12$ $5.21$ $29.89$ $88.28$ $64.22$ $46.78$ $82.15$
@@ A.4 Analyzing Embeddings
Figure 7: An overview flowchart of all datasets used in the experiments of this paper.
@@ Appendix B Annotations Statistics
Table 8: Annotation statistics for 100 doctor-written (Hallucinations-MIMIC-DI) and 100 generated (Hallucinations-Generated-DI) patient summaries. Annotation were done separately and agreed on together.
@@ Appendix C Parameter Tuning
[…]
MIMIC-IV-Note-Ext-DI-BHC (100,175 examples)
LED-base (80,140 ex.) $43.32$ $17.05$ $8.26$ $4.30$ $29.21$ $87.98$ $63.52$ $46.39$ $74.36$
LED-large (80,140 ex.) $43.82$ $17.33$ $8.85$ $4.92$ $29.89$ $88.11$ $64.12$ $46.71$ $76.99$
Llama 2 7B (100 ex.) $38.36$ $12.66$ $5.13$ $2.24$ $24.73$ $85.68$ $60.23$ $44.12$ $73.13$
Llama 2 70B (100 ex.) $40.58$ $14.31$ $6.09$ $2.74$ $26.19$ $86.30$ $61.89$ $45.16$ $76.90$
GPT-4 5-shot (5 ex.) $38.80$ $10.78$ $3.55$ $1.12$ $21.98$ $86.67$ $61.30$ $42.88$ $131.86$
@@
MIMIC-IV-Note-Ext-DI-BHC-Anno (26,178 examples)
LED-base (20,942 ex.) $42.30$ $14.98$ $7.04$ $3.87$ $26.50$ $86.71$ $60.85$ $44.38$ $117.81$
LED-large (20,942 ex.) $46.21$ $17.38$ $8.72$ $5.14$ $28.87$ $87.50$ $63.52$ $45.84$ $117.59$
Llama 2 7B (100 ex.) $36.95$ $11.92$ $5.12$ $2.53$ $22.73$ $82.44$ $57.07$ $42.41$ $100.54$
Llama 2 70B (100 ex.) $41.82$ $13.63$ $5.77$ $2.66$ $24.83$ $86.43$ $61.34$ $43.86$ $114.08$
GPT-4 5-shot (5 ex.) $43.07$ $12.91$ $4.79$ $2.13$ $23.91$ $86.80$ $62.08$ $43.40$ $159.68$
@@
Table 11: All performance results of patient summary generation. We tested all models on the full MIMIC-IV-Note-Ext-DI-BHC dataset and the subset MIMIC-IV-Note-Ext-DI-BHC-Anno used for labeling. Parameter or prompt tuning was performed for all models as described in Appendix C.
We performed parameter tuning for the LED (Beltagy et al., 2020) and the Llama 2 (Touvron et al., 2023) models. We used the full MIMIC-IV-Note-Ext-DI-BHC containing 100,175 examples and the subset MIMIC-IV-Note-Ext-DI-BHC-Anno of 26,178 examples used for labeling. We performed a separate parameter tuning on each dataset. For the LED model we performed full fine-tuning using 80% of the data for training. For Llama 2 we used LoRA for parameter-efficient fine-tuning (Hu et al., 2021) on 100 training examples and loaded the model in 8 bit. For both models, we used 100 examples for validation, and 100 examples for testing. The smaller number of validation and testing examples were chosen to reduce the runtime, however, we still consider them sufficient for representative results. We tracked our experiments with Weights & Biases (Biewald, 2020).
@@ Appendix D Prompt Tuning GPT-4
This finding was supported by the quantitative analysis on ten held-out examples from the MIMIC-IV-Note-Ext-DI dataset. The results are shown in table 12. In general, prompt three lead to the best results and additional in-context example did not degrade the performance. Guiding the output with the instructions to start with “You were admitted” also lead to a quantitative improvement. Hence, we chose this prompt for all our GPT-4 summarization experiments.
@@ Appendix E Hallucination Detection with GPT-4
Table 13: Recall for different hallucination labels for hallucination detection on 100 doctor-written summaries (Hallucinations-MIMIC-DI) and 100 generated summaries (Hallucinations-Generated-DI) using partial matching.
@@
Table 15,16,17 show the prompt used to inform GPT-4 of potential error types and the format for annotating the hallucination spans. We can optionally turn off the hallucination class detection (i.e., generating only <error> rather than <error class="hallucination type">), and we compare the results as class-aware and class-agnostic hallucination detection.
One can also optionally prompt the LLM to reason about potential spans of hallucination before annotating the full text with chain-of-thought reasoning. In practice, we prompt GPT-4 to generate a list of labeled hallucination and explanations that mimics the examples in the prompt, i.e., generating bullet point lists like - "Your <error>red blood cell count</error> was followed and was stable." The BHC does not state that the red blood cell count was followed. Instead the hematocrit remained stable according to the BHC. with explanation before annotating the full summary with hallucination labels.
1. Relevance: The rating measures how well the summary captures the key points of the brief hospital course. Consider whether all and only the important aspects are contained in the summary. To this end, please label key facts in the context and the summary.
5 = all key points included
4 = 1 key point missing
3 = 2 key point missing
2 = 3 key point missing
1 = more than 3 key points missing
2. Consistency: The rating measures whether the facts in the summary are consistent with the facts in the original brief hospital course. Consider whether the summary does reproduce all facts accurately and does not make up untrue information. To this end, please label all hallucination in the summary according to the protocol.
5 = no finding
4 = 1-2 minor findings
3 = more than 2 minor or 1 major finding
2 = 2 major findings
1 = more than 2 major findings
3. Simplification: The rating measures whether the summary is written in plain language understandable for a patient. Consider medical terms and abbreviations. To this end, please label medical jargon in the summary.
5 = no medical jargon
4 = 1-2 minor unexplained medical term
3 = more than 2 minor or 1 major medical term
2 = 2 major medical terms
1 = more than 2 major medical terms
4. Fluency: This rating measures the quality of individual sentences, are they well-written and grammatically correct. Consider the quality of individual sentences.
5 = all sentences are well-written and correct
4 = 1-2 sentences have minor errors or poor style
3 = more than 2 sentences with minor errors or poor style or 1 sentence with severe error
2 = 2 sentences with major errors
1 = more than 3 major errors
5. Coherence: The rating measures the quality of all sentences collectively, to the fit together and sound naturally. Consider the quality of the summary as a whole.
5 = the summary as a whole is well-written and clearly structured
4 = the summary is mostly well-written and structured
3 = the summary can be followed and has some structure
2 = the summary is hard to follow and has no clear structure
1 = the summary is very hard to follow and very unstructured
@@ F.1 Labeling Key Facts and Medical Jargon
Figure 10: Example for the creation of the derived datasets Cleaned and Cleaned & Improved. The example is a slightly modified version of Figure 2. For Cleaned, we manually replaced hallucinations with useful facts in the context or removed them otherwise. For Cleaned & Improved, we further corrected mistakes and artifacts such as typos or deidentifications. We performed this process for all 100 examples in the Original dataset.
@@
User You will be given a doctor’s note and you will need to summarize the patient’s brief hospital course. Let’s do a practice round.
@@ 3. Incorrect fact
See pages 1 of instructions.pdf See pages 2-6 of instructions.pdf